OpenStudy (anonymous):

I need help with these substitution problems..and how to graph them! MEDALS WILL BE GIVEN 1) y=4x-6 when y=2x-4 (x ,y ) 2) y+2x=1 when y=3x+6 (x ,y ) 3) y=x+10 when y=-2x (x ,y ) 4) y= -x+9 when y= -3x (x ,y )

OpenStudy (anonymous):

A solution could not be calculated.

OpenStudy (anonymous):

where they both meet, they can

OpenStudy (anonymous):

@freakenpurplepizza see? It can be calculated.. I just understand this!

OpenStudy (ranga):

I will do the first one, you can follow the same procedure for the rest: 1) y = 4x - 6 (1) y = 2x - 4 (2) Replace y with 4x - 6 in (2) 4x - 6 = 2x - 4 4x - 2x = 6 - 4 2x = 2 x = 1 Put x = 1 in (1) y = 4(1) - 6 = -2 So (1,-2) is the solution. For graphing y = 4x - 6, find where the graph cuts the x and y axes (that is, the x and y intercepts): To find the x-intercept, set y = 0 and solve for x: 0 = 4x - 6 4x = 6 x = 1.5 To find the y-intercept, set x = 0 and solve for y: y = 4(0) - 6 = -6 Mark off 1.5 on the x-axis and -6 on the y-axis and draw a line passing through the two pints using a ruler.

OpenStudy (anonymous):

That makes sense! Alrighty I will follow this procedure and will read this over.. Thank you

OpenStudy (ranga):

You are welcome.